OpenStudy (anonymous):

Consider the equation x^2=e^2+2. Sarting with x1=-1, apply interation of new tons method to compute approximate solutions x2

OpenStudy (anonymous):

so all i have to do is to derivative the equation?

OpenStudy (anonymous):

then x1-(f(x)/(f(x)

OpenStudy (anonymous):

it is really \[x^2=e^2+2\]?

OpenStudy (anonymous):

No sorry, its x^2=e^x+2

OpenStudy (anonymous):

that is what i though start with \[f(x)=e^x-x^2+2\] and then \(f'(x)=e^x-2x\)

OpenStudy (anonymous):

then \[x_2=-1-\frac{f(-1)}{f'(-1)}\] and probably a calculator

OpenStudy (anonymous):

did you drivative?

sam (.sam.):

Hmm I don't see the need of using integration here, you can straight away approximate x from \[\large x^2=e^x+3 \\ \\ \large x_{n+1}=\pm \sqrt{e^{x_n}+3}\] Then use -1 for \(+\sqrt{...} ~~~and~~~-\sqrt{...}\) See if which gives you a constant value throughout the calcs

OpenStudy (anonymous):

so i do not need to drivitative the equation?

sam (.sam.):

Yeah try it out you should get some recurring numbers

OpenStudy (anonymous):

sorry the proper equation is x^2=e^x+2 @.Sam.

sam (.sam.):

Oh doesn't matter, change the 3 to 2

sam (.sam.):

\[\large x^2=e^x+2 \\ \\ \large x_{n+1}=\pm \sqrt{e^{x_n}+2}\] Iterate when x=-1